OpenStudy (anonymous):

please HELP!!!! MEDAL & FAN. If f(x) = x + 4 and g(x) = 2x - 3, find (g - f)(2)

hero (hero):

(g - f)(x) = g(x) - f(x): Therefore: g(x) - f(x) = 2x - 3 - (x + 4) = 2x - 3 - x - 4 = (2x - x) - (3 + 4) = x - 7 Thus: (g - f)(2) = g(2) - f(2) = 2 - 7 = - 5

OpenStudy (anonymous):

so what do I do next

hero (hero):

I posted the complete steps bro. Those are ALL the steps.

OpenStudy (anonymous):

that's the answer ?

OpenStudy (anonymous):

or do I solve

hero (hero):

What I posted were the solution steps. The solution is included.

OpenStudy (anonymous):

so its -5 ?

hero (hero):

Yes, correct.